# Quadrupole-hexadecapole correlations in neutron-rich samarium and gadolinium isotopes

### v1

@@ I Introduction
The importance of studying deformations in nuclei and their effects on various nuclear properties, e.g. excitation energies and transition strengths, has been recognized for decades [1, 2]. The most commonly present and extensively studied nuclear deformation is of quadrupole type, and their effects on the positive-parity states are well known. The higher-order correlations which affect the positive-parity states, the hexadecapole correlations, have been significantly less studied, due to their effects often being overshadowed by large quadrupole correlations. Nevertheless, the presence of hexadecapole correlations has been found in a wide range of nuclei [3, 4], and the main effect of such correlations is the appearance of the $4^{+}$ band in the low-lying excitation energy spectra of some even-even nuclei, with an enhanced $B(E4)$ transition strength from the $4^{+}$ band head state to the $0^{+}$ ground state. Hexadecapole correlations have also been found to be important in describing certain nuclear reactions and decays. [4, 5, 6].
@@
Recently, Gogny Hartree-Fock-Bogoliubov (HFB) calculations were performed to compute potential energy surfaces (PESs) have been computed with constraints on the quadrupole and hexadecapole deformation parameters [34]. In this study, we carry out the mapping procedure by using the Gogny-D1S PESs in order to examine the effects of the quadrupole-hexadecapole coupling on the low-lying spectra of the considered nuclei. Since in previous studies [22, 23], calculations were performed within the framework of the constrained relativistic mean field (MDC-RMF) model [35, 36], with the relativistic density-dependent point-coupling (DD-PC1) functional [37, 38], this work will allow us to study how the quadrupole-hexadecapole mapping method depends on the choice of the EDF. On the other hand, we analyze positive parity bands in heavy Sm and Gd isotopes as well as the effects of hexadecapole collectivity in heavy neutron-rich nuclei, far from the valley of stability. The results of the calculations are compared with the ones obtained with a simpler $sd$-IBM, as well as with the available experimental data [39].
The paper is organized as follows. The theoretical framework, i.e., the Gogny-D1S HFB approach and the $sd$-IBM and $sdg$-IBM, is outlined in Sec. II. Potential energy surfaces computed with both SCMF and IBM calculations are presented in Section III. Spectroscopic properties, including the excitation spectra of low-lying states, electric quadrupole, hexadecapole, and monopole transition properties, are discussed in Sec. IV. Section V is devoted to the concluding remarks and work perspectives.
@@ II Model description
In order to obtain the SCMF PES for the studied nuclei, the HFB equations have been solved with constraints on the axial quadrupole $\hat{Q}_{20}$ and hexadecapole $\hat{Q}_{40}$ moment operators [34]. The axial quadrupole $\beta_{20}\equiv\beta_{2}$ and hexadecapole $\beta_{40}\equiv\beta_{4}$ deformation parameters are related to the expectation values of the multipole moment operators as [34]:
@@
In order to calculate low-lying excitation energies and transition strengths, we have employed the $sdg$-IBM-1 model, referred to simply as $sdg$-IBM. The Hamiltonian is built in the same way as in Ref. [23]:
@@
The parameters $\epsilon_{d}$, $\epsilon_{g}$, $\kappa$, $\chi$ and $\sigma$ are determined by the mapping procedure [16]. The $sdg$-IBM is connected to the geometric model through a coherent state [43], |$\ket{\phi}\sim(1+\tilde{\beta}_{2}d^{\dagger}_{0}+\tilde{\beta}_{4}g^{\dagger}_{0})^{N_{B}}\ket{0}$, with $N_{B}$ representing the number of bosons, determined by the nearest doubly-magic nucleus, and |$\ket{0}$ representing the boson vacuum. Note that $\tilde{\beta}_{2}$ and $\tilde{\beta}_{4}$ represent the axial bosonic quadrupole and hexadecapole deformation parameters, respectively. For Gd and Sm isotopes with $N<106$, the core is represented by the doubly magic nucleus 132Sn. In the case of 168Sm and 170Gd, the corresponding core is 176Sn since the neutrons occupy the upper half of the neutron major shell $N=82-126$.
@@
where the $sd$-IBM quadrupole operator reads:
@@
with $\hat{Q}^{(2)}_{sdg/sd}$ representing the quadrupole operator of the $sdg$ (3) or $sd$ (7) boson model. The hexadecapole operator is chosen, in the case of $sdg$-IBM, to be:
@@
The quadrupole and hexadecapole effective charges $e_{2}^{sdg/sd},e_{4}^{sdg/sd}$ are fitted in order to reproduce the measured transition strengths $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ and $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$, from the first $2^{+}$ and $4^{+}$ states to the ground state $0^{+}$, respectively. The monopole transition operator [44] reads
@@
Figure 1: Axially-symmetric quadrupole ($\beta_{20}$) and hexadecapole ($\beta_{40}$) constrained potential energy surfaces for 150-168Sm. The energy difference between neighbouring contours is 0.1 MeV. The absolute minimum is indicated by an open triangle. Results have been obtained with the Gogny-D1S energy density functional.
Figure 2: The same as in Fig. 1, but for 152-170Gd.
Figure 3: The same as in Fig. 1, but for the mapped $sdg$-IBM potential energy surfaces of 150-168Sm.
Figure 4: The same as in Fig. 1, but for the mapped $sdg$-IBM potential energy surfaces of 152-170Gd.
@@ III Mapping the SCMF results onto the IBM space
The corresponding (mapped) $sdg$-IBM PESs are depicted in Figs. 3 and 4. The mapping procedure reproduces basic features of the fermionic PESs such as the position of the absolute minima, the oblate saddle points for $N=88$ nuclei as well as the overall shape of the surface. The strong quadrupole-hexadecapole coupling is well reproduced in $N\leq 100$ isotopes, while in $N>100$ isotopes, the coupling is predicted to be less strong. Note that the IBM PESs are significantly flatter than the HFB ones, due to the restricted boson (valence) space as compared to the Gogny-HFB model space. This also explains why the mapping procedure is not able to reproduce the “elongated” shapes of the PESs for $N\geq 100$ nuclei. The 1D $sd$-IBM mapping has also been carried out in all the studied isotopes to reproduce some of the properties of the $E_{\textnormal{SCMF}}(\beta_{2},\beta_{4}=0)$ curve such as the position of the absolute minimum, the position and energy of the oblate saddle point, and the energy of the spherical $\beta_{2}=0$ configuration.
@@
The $sdg$- and $sd$-IBM parameters are shown in Figs. 5 and 6, as functions of the neutron number $N$. The $d$-boson energy $\epsilon_{d}$ has a very similar behaviour in both models. It decreases up to $N=100$, after which it stops decreasing, and even increases at $N=106$, which corresponds to lower $\beta_{2}^{\rm min}$ values. Moreover, the parameters $\kappa$, $\chi$ and $C_{2}$ also display a similar behaviour in both models, whereas in the $sdg$-IBM the $\chi$ parameter is characterized by significantly lower absolute values, compared to the $sd$-IBM ones. Regarding those parameters appearing only in the $sdg$-IBM, the $g$-boson energy $\epsilon_{g}$ takes values between $\epsilon_{g}=1.1$ and $\epsilon_{g}=1.3$ MeV. The $C_{4}$ parameter decreases until $N=96$, after which it starts to increase, which corresponds to $\beta_{4}^{\rm min}$ values becoming lower for $N\geq 96$ isotopes. Finally, with increasing neutron number, the parameter $\sigma$, which accounts for the quadrupole-hexadecapole coupling in the $sdg$-IBM, drops from $\sigma=2.8$ to $\sigma=1.0$, which also corresponds to a decrease in $\beta_{4}^{\textnormal{min}}$. This explains why the coupling is predicted to be less strong in $N>100$ isotopes by the $sdg$-IBM. It should also be noted that $\sigma$ decreases faster in Gd isotopes, compared to the Sm ones, which can be explained from the fact that $\beta_{4}^{\textnormal{min}}$ values are smaller in Gd isotopes with the same $N$.
@@ IV.1 Excitation energies
Figure 7 shows the excitation energies corresponding to the yrast band states with even spin $J^{\pi}=2^{+}$ to $14^{+}$. As expected, the $sdg$-IBM improves the description of the excitation energies of higher-spin states with $J^{\pi}=12^{+}$ and $14^{+}$ in transitional $N=88$ nuclei. For Sm nuclei, the $sdg$-IBM gives a better description of the energy levels up to $N=92$ than the $sd$-IBM. This can be explained by the fact that large amounts of $g$-boson contributions are present in the higher-spin states, as indeed, the expectation value of the $g$-boson number operator $\braket{\hat{n}_{g}}$ is greater or equal to 1 in those isotopes with $N\leq 92$. Note that for Sm isotopes with $N\leq 92$, the $\beta_{4}^{\rm min}$ values are larger compared to those for Gd isotopes with the same $N$. For the $N>92$ isotopes, there is no significant difference between both models concerning the description of the ground state bands. This is not surprising, since as the number of valence nucleons increases for deformed nuclei even the $sd$-IBM is good enough to provide an accurate description of the yrast-band levels. In any case, in general the $sdg$-IBM does provide some significant improvements in the description of the yrast band, compared to the simpler $sd$-IBM.
@@
Figure 9 shows the excitation energies of the $2^{+}_{2}$, $3^{+}_{1}$, and $4^{+}_{2}$ states, which can be identified as members of the $\gamma$-vibrational band in the present calculation. Both models predict similar energies for Sm and Gd isotopes. Those energies mostly agree well with the experimental data. The description of the $\gamma$-vibrational band could be improved by including both three-body terms [19] and the $\gamma$ degree of freedom at the HFB level. However, such an extension is beyond the scope of this paper. Work along these lines will be reported in future publications.
@@ IV.2.1 $E2$ transitions
Figure 10 shows the $B(E2;J\rightarrow J-2)$ transition strengths in the ground state bands of the well deformed $N=90$ and 92 isotones. These nuclei are specifically considered, since there are experimental data on $E2$ transitions. In Sm isotopes, the $sdg$-IBM significantly improves the description of $B(E2)$ transition strengths for $J^{\pi}\geq 6^{+}$. However, for 154Gd, the $sdg$-IBM predicts $B(E2)$ strengths significantly larger than the $sd$-IBM, which are however still smaller than the measured values. The discrepancy points towards some missing correlations other than the hexadecapole ones, that could also contribute to the $B(E2)$ strengths. Due to the large experimental error bars, in 156Gd it is not possible to conclude whether $sdg$-IBM improves the description of the $E2$ transitions in the yrast band or not. The fact that the $sdg$-IBM predicts larger $B(E2)$ transition strengths than the $sd$-IBM can be attributed to the large values of the parameter $\sigma$ (see Fig. 5). In the case of 152,154Sm and 154Gd the value $\sigma=2.8$ leads to a significant contribution to the $B(E2)$ transition strengths arising, from the $\left[d^{\dagger}\times\tilde{g}+g^{\dagger}\times\tilde{d}\right]^{(2)}$ part of the quadrupole $\hat{Q}^{(2)}$ operator of Eq.(3). Moreover, the $\sigma=2.1$ value obtained for 156Gd leads to similar $sdg$-IBM and $sd$-IBM $B(E2)$ strengths. Note that the $\sigma$ values for 154,156Gd are larger than those previously obtained [23] using microscopic input from the relativistic DD-PC1 EDF.
@@ IV.2.2 Hexadecapole transitions
The $B(E4;4_{n}^{+}\rightarrow 0_{1}^{+})$ ($n=1,2,3,4$) reduced transition probabilities obtained in the mapped calculations are plotted in Fig. 11, as functions of the nucleon number $A$ in order to avoid overlapping between the data for Sm and Gd isotopes. The $e_{4}^{sdg/sd}$ effective charges are fitted to the experimental $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$ values [39, 46, 47, 48]. Results for $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$ strengths are plotted in panels (a) and (b) of the figure, while $B(E4)$ values from non-yrast $4_{n}^{+}$ (with $n=2,3,4$) to the $0^{+}$ ground states are shown in panels (c)-(h). The $4_{2}^{+}$ state typically belongs to the $\gamma$-vibrational band, where $g$ bosons do not play a significant role. Thus both models predict weak $E4$ transitions in this case. Large $sdg$-IBM $B(E4)$ values are obtained for the $4_{3}^{+}$ states in 150Sm and 152Gd. The strong $E4$ transition, predicted for the $4_{4}^{+}$ state in 154Gd, agrees well with the experimental identification of this state as a band head of the $K=4^{+}$ band [39]. On the other hand, the $sd$-IBM gives negligibly small $B(E4)$ reduced transition probabilities. This indicates a significant difference between the $sdg$-IBM and the simpler $sd$-IBM. Nevertheless, experimental data on $B(E4)$ strengths in rare-earth nuclei are still required to understand whether the $sdg$-IBM describes them better.

### v2

@@ I Introduction
The importance of studying deformations in nuclei and their effects on various nuclear properties, e.g. excitation energies and transition strengths, has been recognized for decades [1, 2]. The most commonly present and extensively studied nuclear deformation is of quadrupole type, and its effects on the positive-parity states are well known. The higher-order correlations which affect the positive-parity states, the hexadecapole correlations, have been significantly less studied, due to their effects often being overshadowed by large quadrupole correlations. Nevertheless, the presence of hexadecapole correlations has been found in a wide range of nuclei [3, 4], and the main effect of such correlations is the appearance of the $4^{+}$ band in the low-lying excitation energy spectra of some even-even nuclei, with an enhanced $B(E4)$ transition strength from the $4^{+}$ bandhead state to the $0^{+}$ ground state. Hexadecapole correlations have also been found to be important in describing certain nuclear reactions and decays [4, 5, 6].
@@
Recently, Gogny Hartree-Fock-Bogoliubov (HFB) calculations, with constraints set on the quadrupole and hexadecapole deformation parameters, were performed to compute potential energy surfaces (PESs) in rare-earth nuclei [34]. In this study, we carry out the mapping procedure by using the Gogny-D1S PESs in order to examine the effects of the quadrupole-hexadecapole coupling on the low-lying spectra of the considered nuclei. Since, in previous studies [22, 23], calculations were performed within the framework of the constrained relativistic mean field (MDC-RMF) model [35, 36], with the relativistic density-dependent point-coupling (DD-PC1) functional [37, 38], this work will allow us to study how the quadrupole-hexadecapole mapping method depends on the choice of the EDF. On the other hand, we analyze positive-parity bands in heavy Sm and Gd isotopes as well as the effects of hexadecapole collectivity in heavy neutron-rich nuclei, far from the valley of stability. The results of the calculations are compared with the ones obtained with a simpler $sd$-IBM, as well as with the available experimental data [39].
The paper is organized as follows. The theoretical framework, i.e., the Gogny-D1S HFB approach and the $sd$-IBM and $sdg$-IBM, is outlined in Sec. II. Potential energy surfaces computed with both SCMF and IBM calculations are presented in Sec. III. Spectroscopic properties, including the excitation spectra of low-lying states, electric quadrupole, hexadecapole, and monopole transition properties, are discussed in Sec. IV. Section V is devoted to the concluding remarks and work perspectives.
@@ II Model description
In order to obtain the SCMF PES for the studied nuclei, the HFB equations have been solved with constraints on the axial quadrupole $\hat{Q}_{20}$ and hexadecapole $\hat{Q}_{40}$ moment operators [34]. The axial quadrupole $\beta_{20}\equiv\beta_{2}$ and hexadecapole $\beta_{40}\equiv\beta_{4}$ deformation parameters are related to the expectation values of the multipole moment operators as [34]
@@
In order to calculate low-lying excitation energies and transition strengths, we have employed the $sdg$-IBM-1, referred to simply as $sdg$-IBM. The Hamiltonian is built in the same way as in Ref. [23]:
@@
The parameters $\epsilon_{d}$, $\epsilon_{g}$, $\kappa$, $\chi$ and $\sigma$ are determined by the mapping procedure [16]. The $sdg$-IBM is connected to the geometric model through a coherent state [43], |$\ket{\phi}\sim(1+\tilde{\beta}_{2}d^{\dagger}_{0}+\tilde{\beta}_{4}g^{\dagger}_{0})^{N_{B}}\ket{0}$, with $N_{B}$ representing the number of bosons, determined by the nearest doubly magic nucleus, and |$\ket{0}$ representing the boson vacuum. Note that $\tilde{\beta}_{2}$ and $\tilde{\beta}_{4}$ represent the axial bosonic quadrupole and hexadecapole deformation parameters, respectively. For Gd and Sm isotopes with $N<106$, the core is represented by the doubly magic nucleus 132Sn. In the case of 168Sm and 170Gd, the corresponding core is 176Sn since the neutrons occupy the upper half of the neutron major shell $N=82-126$.
@@
where the $sd$-IBM quadrupole operator reads
@@
with $\hat{Q}^{(2)}_{sdg/sd}$ representing the quadrupole operator of the $sdg$ (3) or $sd$ (7) boson model. The hexadecapole operator is chosen, in the case of $sdg$-IBM, to be
@@
The quadrupole and hexadecapole effective charges $e_{2}^{sdg/sd}$ and $e_{4}^{sdg/sd}$ are fitted in order to reproduce the measured transition strengths $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ and $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$, from the first $2^{+}$ and $4^{+}$ states to the ground state $0^{+}$, respectively. The monopole transition operator [44] reads
@@
Figure 1: Axially symmetric quadrupole ($\beta_{20}$) and hexadecapole ($\beta_{40}$) constrained potential energy surfaces for 150-166Sm. The energy difference between neighbouring contours is 0.1 MeV. The absolute minimum is indicated by an open triangle. Results have been obtained with the Gogny-D1S energy density functional.
Figure 2: The same as in Fig. 1, but for 152-168Gd.
Figure 3: The same as in Fig. 1, but for the mapped $sdg$-IBM potential energy surfaces of 150-166Sm.
Figure 4: The same as in Fig. 1, but for the mapped $sdg$-IBM potential energy surfaces of 152-168Gd.
@@ III Mapping the SCMF results onto the IBM space
The corresponding (mapped) $sdg$-IBM PESs are depicted in Figs. 3 and 4. The mapping procedure reproduces basic features of the fermionic PESs such as the position of the absolute minima, the oblate saddle points for $N=88$ nuclei as well as the overall shape of the surface. The strong quadrupole-hexadecapole coupling is well reproduced in $N\leq 100$ isotopes, whereas in $N>100$ isotopes, the coupling is predicted to be less strong. Note that the IBM PESs are significantly flatter than the HFB ones, due to the restricted boson (valence) space as compared to the Gogny-HFB model space. This also explains why the mapping procedure is not able to reproduce the “elongated” shapes of the PESs for $N\geq 100$ nuclei. The 1D $sd$-IBM mapping has also been carried out in all the studied isotopes to reproduce some of the properties of the $E_{\textnormal{SCMF}}(\beta_{2},\beta_{4}=0)$ curve such as the position of the absolute minimum, the position and energy of the oblate saddle point, and the energy of the spherical $\beta_{2}=0$ configuration.
@@
The $sdg$- and $sd$-IBM parameters are shown in Figs. 5 and 6, as functions of the neutron number $N$. The $d$-boson energy $\epsilon_{d}$ has a very similar behavior in both models. It decreases up to $N=100$, after which it stops decreasing, and even increases at $N=106$, which corresponds to lower $\beta_{2}^{\rm min}$ values. Moreover, the parameters $\kappa$, $\chi$ and $C_{2}$ also display a similar behavior in both models, whereas in the $sdg$-IBM the $\chi$ parameter is characterized by significantly lower absolute values, compared to the $sd$-IBM ones. Regarding those parameters appearing only in the $sdg$-IBM, the $g$-boson energy $\epsilon_{g}$ takes values between $\epsilon_{g}=1.1$ and $\epsilon_{g}=1.3$ MeV. The $C_{4}$ parameter decreases until $N=96$, after which it starts to increase, which corresponds to $\beta_{4}^{\rm min}$ values becoming lower for $N\geq 96$ isotopes. Finally, with increasing neutron number, the parameter $\sigma$, which accounts for the quadrupole-hexadecapole coupling in the $sdg$-IBM, drops from $\sigma=2.8$ to $\sigma=1.0$, which also corresponds to a decrease in $\beta_{4}^{\textnormal{min}}$. This explains why the coupling is predicted to be less strong in $N>100$ isotopes by the $sdg$-IBM. It should also be noted that $\sigma$ decreases faster in Gd isotopes, compared to the Sm ones, which can be explained from the fact that $\beta_{4}^{\textnormal{min}}$ values are smaller in Gd isotopes with the same $N$.
@@ IV.1 Excitation energies
Figure 7 shows the excitation energies corresponding to the yrast band states with even spin $J^{\pi}=2^{+}$ to $14^{+}$. As expected, the $sdg$-IBM improves the description of the excitation energies of higher-spin states with $J^{\pi}=12^{+}$ and $14^{+}$ in transitional $N=88$ nuclei. For Sm nuclei, the $sdg$-IBM gives a better description of the energy levels up to $N=92$ than the $sd$-IBM. This can be explained by the fact that large amounts of $g$-boson contributions are present in the higher-spin states, as indeed, the expectation value of the $g$-boson number operator $\braket{\hat{n}_{g}}$ is greater than or equal to 1 in those isotopes with $N\leq 92$. Note that for Sm isotopes with $N\leq 92$, the $\beta_{4}^{\rm min}$ values are larger compared to those for Gd isotopes with the same $N$. For the $N>92$ isotopes, there is no significant difference between both models concerning the description of the ground state bands. This is not surprising, since as the number of valence nucleons increases for deformed nuclei even the $sd$-IBM is good enough to provide an accurate description of the yrast-band levels. In any case, in general the $sdg$-IBM does provide some significant improvements in the description of the yrast band, compared to the simpler $sd$-IBM.
@@
Figure 9 shows the excitation energies of the $2^{+}_{2}$, $3^{+}_{1}$, and $4^{+}_{2}$ states, which can be identified as members of the $\gamma$-vibrational band in the present calculation. Both models predict similar energies for Sm and Gd isotopes. While the calculated energies of the $\gamma$-vibrational band states are in good agreement with the experimental data for the nuclei with $N\leq 90$ and $N\geq 96$, both models underestimate the energies for those with $N=92$ and 94. Overall, the presence of $g$ bosons in the IBM calculations does not seem to have any significant effects on the $\gamma$-vibrational band. The description of the $\gamma$-vibrational band could be improved by including both three-body terms [19] and the $\gamma$ degree of freedom at the HFB level. However, such an extension is beyond the scope of this paper. Work along these lines will be reported in future publications.
@@ IV.2.1 $E2$ transitions
Figure 10 shows the $B(E2;J\rightarrow J-2)$ transition strengths in the ground state bands of the well deformed $N=90$ and 92 isotones. These nuclei are specifically considered, since there are experimental data on $E2$ transitions. In Sm isotopes, the $sdg$-IBM significantly improves the description of $B(E2)$ transition strengths for $J^{\pi}\geq 6^{+}$. However, for 154Gd, the $sdg$-IBM predicts $B(E2)$ strengths significantly larger than the $sd$-IBM, which are however still smaller than the measured values. The discrepancy points towards some missing correlations other than the hexadecapole ones, that could also contribute to the $B(E2)$ strengths. Due to the large experimental error bars, in 156Gd it is not possible to conclude whether $sdg$-IBM improves the description of the $E2$ transitions in the yrast band or not. The fact that the $sdg$-IBM predicts larger $B(E2)$ transition strengths than the $sd$-IBM can be attributed to the large values of the parameter $\sigma$ (see Fig. 5). In the case of 152,154Sm and 154Gd the value $\sigma=2.8$ leads to a significant contribution to the $B(E2)$ transition strengths arising from the $\left[d^{\dagger}\times\tilde{g}+g^{\dagger}\times\tilde{d}\right]^{(2)}$ part of the quadrupole $\hat{Q}^{(2)}$ operator of Eq.(3). Moreover, the $\sigma=2.1$ value obtained for 156Gd leads to similar $sdg$-IBM and $sd$-IBM $B(E2)$ strengths. Note that the $\sigma$ values for 154,156Gd are larger than those previously obtained [23] using microscopic input from the relativistic DD-PC1 EDF.
@@ IV.2.2 Hexadecapole transitions
The $B(E4;4_{n}^{+}\rightarrow 0_{1}^{+})$ ($n=1,2,3,4$) reduced transition probabilities obtained in the mapped calculations are plotted in Fig. 11, as functions of the nucleon number $A$ in order to avoid overlapping between the data for Sm and Gd isotopes. The $e_{4}^{sdg/sd}$ effective charges are fitted to the experimental $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$ values [39, 46, 47, 48]. Results for $B(E4;4_{1}^{+}\rightarrow 0_{1}^{+})$ strengths are plotted in Fig. 11(a) and 11(b), while $B(E4)$ values from non-yrast $4_{n}^{+}$ (with $n=2,3,4$) to the $0^{+}$ ground states are shown in Figs. 11(c)-11(h). The $4_{2}^{+}$ state typically belongs to the $\gamma$-vibrational band, where $g$ bosons do not play a significant role. Thus both models predict weak $E4$ transitions in this case. Large $sdg$-IBM $B(E4)$ values are obtained for the $4_{3}^{+}$ states in 150Sm and 152Gd. The strong $E4$ transition, predicted for the $4_{4}^{+}$ state in 154Gd, agrees well with the experimental identification of this state as a bandhead of the $K=4^{+}$ band [39]. On the other hand, the $sd$-IBM gives negligibly small $B(E4)$ reduced transition probabilities. This indicates a significant difference between the $sdg$-IBM and the simpler $sd$-IBM. Nevertheless, experimental data on $B(E4)$ strengths in rare-earth nuclei are still required to understand whether the $sdg$-IBM describes them better.